OpenStudy (imsleeziboii):

EASY MEDAL I only have 5 min.

OpenStudy (studygurl14):

What's your question?

OpenStudy (tkhunny):

You already wasted one of your minutes by failing to post your question. I suspect you'll waste another by failing to show YOUR work. Let's see!

OpenStudy (imsleeziboii):

Rewrite the radicand as two factors one of which is a perfect cube. \[^{3}\sqrt{250}\]

OpenStudy (anonymous):

try 5³

OpenStudy (imsleeziboii):

75?

OpenStudy (imsleeziboii):

1 minute i have now

OpenStudy (anonymous):

\[5^3 = 125\] \[\sqrt[3]{250} = \sqrt[3]{125 \times 2} = 5 \times \sqrt[3]{2}\]

OpenStudy (imsleeziboii):

thanks so much

OpenStudy (anonymous):

np

OpenStudy (imsleeziboii):

I got a 20% im done...